# Spatial Context Improves the Integration of Text with Remote Sensing for Mapping Environmental Variables

### v1

@@ 2.1 Preliminaries
We consider a dataset composed of NN geolocated samples with the following elements: the location $x_{i}\in\mathbb{R}^{2}$ is a GPS coordinate; $I_{i}\in\mathbb{R}^{3\times h\times w}$ is a remote sensing image of location $x_{i}$ with RGB bands; $T_{i}=\{s_{i,0},...,s_{i,j}\}$ is the set of $j$ text sentences associated with location $x_{i}$; and $y_{i}\in\mathbb{R}^{m}$ is the set of target numerical environmental variables, while $m$ denotes the number of environmental variables to be predicted.
@@
The proposed approach aims at learning a model $f$ that estimates the environmental variables $y_{*}$ for a new location $x_{*}$, as $\hat{y}_{*}=f(p_{*})$
@@ 2.2 Text and image encoding
Given a training location $p_{i}$, a pretrained visual encoder $f_{v}$ encodes the $k+1$ remote sensing images into $k+1$ visual embeddings $(v_{i})_{i=0}^{k}$ with $v_{i}\in\mathbb{R}^{d}$. For each location, a number $j$ of sentences is then selected for encoding. We found empirically that selecting the top-$j$ sentences per location by computing the cosine similarity with the visual embeddings using a pretrained VLM is the most effective way to integrate text information. More details are available in Section 3.3. The selected sentences are then fed to a tokeniser and to a pretrained text encoder $f_{t}$, producing $j(k+1)$ textual embeddings $(t_{i})_{i=0}^{j(k+1)}$ with $t_{i}\in\mathbb{R}^{d}$.
@@ 2.3 Location encoding
We encode the spatial structure of each textual and visual embedding by combining it with a geolocation encoding based on their geographic coordinates (LOC in Fig. 2). We compare several approaches, while the results are discussed in Section 3.3.2:
  • 1.
@@
    Rank: We encode the rank of nearest neighbours, i.e. the k in the kNN, similarly to NLP approaches that encode the position of a token in a sequence. For that, we use a $512$ dimensional 1D sinusoidal encoding, similarly to the Transformer positional encoding (Vaswani et al., 2017). Then, the embeddings are concatenated with the text and visual features and merged through a fully connected (FC) layer.
  • 3.
@@
    Learnable: Similarly to rank above, we encode the rank of nearest neighbours for both text and image embeddings, but with learnable embeddings. The embeddings are initialized with a uniform distribution between $\frac{-1}{h_{dim}^{1/2}},\frac{1}{h_{dim}^{1/2}}]$ and summed with the corresponding vision and text embeddings.
  • 4.
    Coordinates: We encode the geographic coordinates for each location with a 2D sinusoidal multi-scale encoding, similarly to the “grid” approach in Mai et al. (2020). We scale the geographic coordinates to a relative scale in the range $[0,1]$ based on the min-max coordinates of the study area. We encode each coordinate, i.e. North and East, separately with $h_{dim}/2=256$ dimensions. We concatenate them together with their corresponding text or visual embeddings, and finally combine them through a FC layer.
  • 5.
@@
    Distance: We encode the relative position of the nearest neighbours with the Euclidean distance in meters from the origin to the target nearest neighbour with 1D sinusoidal encoding with $h_{dim}=512$ dimensions up to a maximum distance of $10^{\prime}000$ m. We concatenate the distance embeddings with their corresponding text or visual embeddings and combine them through a FC layer.
  • 6.
@@
    Polar: Mai et al. (2020) showed that using polar coordinates outperforms grid-like approaches. We combine the Euclidean distance from the origin to the target nearest neighbour with the angle between the positive x-axis and the vector pointing from the origin to the nearest neighbour. We use sinusoidal functions to encode both the distance and the angle, similarly to the “polar” approach from Mai et al. (2020). We concatenate the distance, the angle and the features embedding together and combine them through a 1-layer MLP.
@@
After encoding the location for all inputs, each text and visual embedding can be considered as an input token without specific order, since the relative location to the origin is preserved, thanks to the location embeddings.
2.4 Multi-modal fusion with the attention-based module
@@
Once all inputs have been encoded into a common embedding space, fusion can be performed. The fusion module is composed of a transformer encoder, a layer-normalisation layer, a global pooling layer on the global classification (CLS) token and a projection layer that leads to the final model predictions. The attention mechanisms in the transformer encoder enable the fusion of the multi-modal tokens. The motivation behind this design is that the attention mechanism enables to weight dynamically spatial neighbours, emphasising the most informative neighbours, while ignoring distant or irrelevant ones, thereby learning a context-aware spatial representation.
@@ 2.5.1 EcoWikiRS
The EcoWikiRS (Zermatten et al., 2025) dataset contains over $90^{\prime}000$ text-image pairs, combining high-resolution RGB aerial imagery from Switzerland with sentences from Wikipedia describing the habitat of a species observed at the location of the images. The species habitat description and aerial images have been paired together by using the geolocation of crowd-sourced species observations from the Global Biodiversity Information Facility (GBIF) [^1] https://www.gbif.org/. A specificity of this dataset is that most images are paired with more than one sentence, since several sentences are usually extracted from each Wikipedia article, depending on the number of species found and the length of the corresponding Wikipedia article. The dataset is split into train, validation and test following a spatial block split, with a block size of roughly $20$ km, and nearest neighbours are sampled only within the same split. This avoids performance inflation due to spatially autocorrelated training and test samples, when using a random split with geolocated observations (Kattenborn et al., 2022).
@@ 2.5.2 SWECO25
We assess the contribution of each modality in the EcoWikiRS dataset by evaluating its capacity to predict environmental variables. For this task, we use the SWECO25 (Külling et al., 2024) dataset, a data cube of more than $5^{\prime}000$ environmental rasters at $25m$ spatial resolution covering Switzerland. For each location in EcoWikiRS, we extract average values covering the surface of the input image ($100\times 100m$) for 103 selected SWECO25 variables. We standardise the numerical values across the training set for zero-mean and unit variance. The Supplementary Section LABEL:sec:sweco_docu provides more details on the selection, preprocessing and individual description of the SWECO25 variables. The 103 variables are grouped into seven thematic groups, following the groups formed in the SWECO25 dataset. We recall here a short description of each group, and more details are available in the Suppl. Table LABEL:tab:metadata_sweco:
  • 1.
@@ 2.6 Experimental setup
During training, random token masking is applied on the visual and textual tokens, with the probability of $30\%$, to reduce overfitting and increase model robustness to varying NN distances. We use the habitat sentences from the EcoWikiRS data, i.e. sentences relative to the habitat section of the article only, since they were shown to perform best in Zermatten et al. (2025). We use the four sentences with the highest cosine similarity to the image in the SkyCLIP latent space, i.e. top-j sentence with $j=4$. If fewer than four sentences are available, sentences are chosen with replacement. If not mentioned otherwise, a polar location encoding is applied. We evaluated the regression of environmental variables with the Pearson correlation coefficient ($R^{2}$).
@@ 3.1 Combining text and images in a spatial context
Fig. 5 illustrates how the number of nearest neighbours affects the accuracy metrics. Performance increases steadily up to $10$ NN, after which it stagnates or even drops for all models. Including distant neighbours yields no meaningful performance gain while increasing computational cost, arguing for a conservative approach to data integration. This behaviour is better understood when looking at the Suppl. Figs  13-13, which illustrates the mean distance to NN, and the spatial autocorrelation between input observations, measured as the feature similarity as a function of distance. Higher similarity between features occurs up to a distance of $1^{\prime}000$ m for visual embeddings, and 700 m for the text ones, on average, which is close to the average distance of the $10^{th}$ NN, which equals 880 m. This suggests that the spatial attention mechanism implicitly learns geospatial statistics and effectively uses them to weight input features.
@@
Fig. 6 confirms this hypothesis and shows the attention for all visual and text tokens as a function of distance, in the test set. The attention score is computed as the logarithm of the maximum attention score across all transformer encoder layers for the best “Text + Images” model and plotted as distance bins of 100 m. Both modalities exhibit their highest attention values at short distances, which then monotonically decrease up to a distance of about $1^{\prime}000$ m. The visual tokens exhibit, overall, a higher attention score compared to the text tokens.
Figure 6: Attention visualisations for the text and image model on the EcoWikiRS test set. Attention values are log-maximum scores taken across all layers and scaled to $[0,1]$, and they are represented as means and standard deviations, split per distance bins of 100m.
@@ 3.2 Specific contribution of text and images for different types of environmental variables
For all the thematic groups, a gain in performance is obtained when adding spatial context, but to a varying degree. The land use/land cover, bioclimatic, edaphic and population groups are associated with globally high $R^{2}$ scores, and received a significant boost when adding spatial context, with a large relative increase in performance from $0.47$ to $0.61$ ($30\%$) for bioclimatic variables and from $0.40$ to $0.54$ ($35\%$) for the edaphic group. The overall highest $R^{2}$ scores are obtained for the Vegetation group, both with ($0.75$) and without ($0.74$) spatial context, while this small relative improvement when adding spatial context suggests that features at the origin are sufficient to model this group.
For the remaining groups, geological and hydrological, the model performance remains low in both scenarios, with only a slight increase with spatial context. These variables are discretised from categorical variables and seem hardly predictable from the RS perspective; adding context information only provides a moderate benefit.
@@ 3.3.1 Integrating text
Fig. 9 (a) illustrates the results for the “Text only” setting. More sentences lead to an increase in the $R^{2}$ correlation coefficient from $0.10$ to $0.15$, as the models better capture relevant context, improving the overall results. Selecting a small number of informative sentences (top-$j$) is more effective than simply increasing the number of sentences for a given location. For instance, selecting the top-4 sentences leads to an $R^{2}$ value of $0.17$ compared to $0.13$ when randomly sampling 4 sentences. When incorporating top-$j$ sentences from NN, the performance climbs to $0.27$. For the “Text + Images” model in Fig. 9 (b), we do not see a significant increase in performance between selecting top-$j$ sentences or random sentences, probably due to the fact that the performance is driven mostly by the visual modality and incorporating one or several sentences does not add complementary information. However, similar to “Text only” setting, integrating text from spatial neighbours leads to the overall best performance.
@@ 3.3.2 Integrating location
In the presence of visual neighbours, adding absolute seems not to provide any advantage with respect to adding no location information at all (None). As coordinates offers a spatial resolution of approximately 1km1~km when compressed into 512dimensions512~dimensions, it is possibly too coarse for the considered task, with the $10^{th}$ NN lying at 1km1~km on average (Suppl. Fig. 13).
We can see better performances with approaches that encode the relative position of NN to the origin location. The distance performs only slightly better than None, since absolute distance fails to separate neighbours located at similar ranges, but in different directions. Polar and rank obtain the best performance together with learnable. Rank and learnable, seem to be effective because they order NN by decreasing order of spatial correlation, thus theoretically of decreasing importance. Polar encoding is more effective at encoding the spatial context than distances: it provides roughly 40m40~m resolution for distances up to 10km10~km, and the angle resolves neighbours located at similar ranges but in different directions. Overall, Polar obtains similar or best performance, but with fewer trainable parameters, motivating our choice for other experiments.
@@ 4 Discussion and Conclusion
Text appears as an emerging modality in geospatial ecology tasks, giving insights into properties hidden from other traditional data sources. However, the fields of application for which text can contribute are not well defined yet, and its combination with other geospatial data sources is limited by its sparse, irregular, and weakly geolocated nature.
This work addresses these issues in two complementary ways. First, we propose a model architecture combining continuous (imagery) and point-based (text) observations: We incorporate all samples in the vicinity of the point of interest, and preserve their spatial structure with a location encoding approach. Second, we provide a quantitative and qualitative study of how text benefits model performance when combined with RS imagery. We evaluate our approach for predicting 103 environmental variables from SWECO25 (Külling et al., 2024) by relying on the EcoWikiRS dataset (Zermatten et al., 2025), composed of aerial imagery and text extracted from Wikipedia, that describes the local environmental context over the Swiss territory.
With respect to the literature, Space2Vec (Mai et al., 2020) is probably the work most similar to our approach, which proposes various spatial encoding methods to classify images associated with their geolocation and their spatial neighbours, and uses an attention-based module to combine the input variables. Similarly to Kriging, Space2Vec integrates target values from the nearest neighbours as predictors in their model, which are not included in our model. Another difference to their approach is that we combine features coming from different modalities (text and images), whereas they focus on tabular data only. Our model also shows some similarities with (Lange et al., 2025), which combines several pairs of geographic coordinates for estimating species ranges, with an attention-based fusion module. The main difference to our work is that they consider geographic coordinates to be the main inputs, and a text, such as species descriptions or pictures are seen as optional metadata.
@@
Our results demonstrated that combining text with RS imagery beats approaches relying only on one individual modality, i.e. ”Text only” or ”Image only” approach. Moreover, the proposed integration of spatial neighbours consistently provides an additional gain in performance, underlining the benefits of incorporating spatial information for modelling geospatial processes. Text is most beneficial for variables governed by broad-scale processes (climate, edaphic, population), while variables driven by local features (vegetation) remain dominated by imagery. Variables that were poorly predicted from remote sensing alone (e.g., geological and hydrological variables) exhibit a marginal gain from spatial context, and no clear benefit from text integration, suggesting that insufficient relevant information was found in the text data.
Despite promising results, this work presents several potential limitations. First, the scale of the study is limited to Switzerland. While some text-image datasets cover the entire Earth (Daroya et al., 2025), no dataset similar to SWECO25 exists at a global scale with a sufficient spatial resolution, to the best knowledge of the authors, limiting the potential to evaluate such an approach at the global scale. Second, our experiments only evaluate exploiting Wikipedia articles in English. Extending to multilingual settings and additional textual resources beyond Wikipedia could improve coverage and robustness. Finally, the risk of spatial overfitting cannot be excluded, particularly whether specific “prototype” words dominate predictions at certain locations and how this affects generalisation. Despite the spatial block split, the model may learn spurious correlations and sampling biases rather than true semantic content, an issue difficult to quantify in experimental settings.
Overall, this work demonstrates that sparse, weakly geolocated text can reliably improve spatial regression of environmental variables when integrated through an explicit spatial context model. We hope this work encourages further integration of knowledge from more diverse resources for various applications at the global scale.
@@
Supplementary Material
4.1 Spatial statistics of the EcoWikiRS dataset
Fig. 13 shows the mean distance to nearest neighbours for samples from the EcoWikiRS train set. Fig. 13 shows the image-to-image and text-to-text cosine similarity as a function of distance for samples from the EcoWikiRS train set, encoded with the SkyCLIP model.
Figure 12: Mean distance to $k^{th}$ nearest neighbours for samples from the EcoWikiRS dataset, with mean and distribution for 5955-95 and 1991-99 percentiles.
Figure 13: Cosine similarity between image features (a) and text features (b) as a function of the distance between the origin and the NN samples and separated into distance lag of $100m$. (c) Number of pairs of samples used to compute the similarity plots above. All values have been computed on the training set of the EcoWikiRS dataset. The horizontal axis is shared among the three plots.
4.2 Additional results
Figure 14: Scatter plots of SWECO25variables and their predicted values. Four variables per thematic group are randomly plotted. The red line represents the $x=y$ line. Predicted values between 1% and 99% percentiles are plotted (excluding outliers for readability reasons).

### v2

@@ 2.1 Preliminaries
We consider a dataset composed of NN geolocated samples with the following elements: the location $x_{i}\in\mathbb{R}^{2}$ is a GPS coordinate; $I_{i}\in\mathbb{R}^{3\times h\times w}$ is a remote sensing image of location $x_{i}$ with RGB bands; $T_{i}=\{s_{i,0},...,s_{i,j}\}$ is the set of $j$ text sentences associated with location $x_{i}$; and $y_{i}\in\mathbb{R}^{m}$ is the regression target, while $m$ denotes the number of environmental variables to be predicted.
@@
The proposed approach aims at learning a model $f$ that estimates the environmental variables $y_{*}$ for a new location $x_{*}$, as $\hat{y}_{*}=f(p_{*})$.
@@ 2.2 Text and image encoding
Given a training location $p_{i}$, a pretrained visual encoder $f_{v}$ encodes the $k+1$ remote sensing images into $k+1$ visual embeddings $(v_{i})_{i=0}^{k}$ with $v_{i}\in\mathbb{R}^{d}$. For each location, a number $j$ of sentences is then selected. The selected sentences are fed to a tokeniser and to a pretrained text encoder $f_{t}$, producing $j(k+1)$ textual embeddings $(t_{i})_{i=0}^{j(k+1)}$ with $t_{i}\in\mathbb{R}^{d}$. We found empirically that selecting the top-$j$ sentences per location by computing the cosine similarity with the visual embeddings using a pretrained VLM is the most effective way to integrate relevant text information. More details are available in Section 3.3.
@@ 2.3 Location encoding
We encode the spatial structure of each textual and visual embedding by combining it with a geolocation encoding based on their geographic coordinates ($LOC_{i}$ in Fig. 2). While several studies propose pretrained location embeddings available at a global scale, we consider them unsuitable for our study area, which is limited to a single country (Switzerland). Such embeddings would likely produce overly homogeneous representations. Following Mai et al. (2020), we compare several learnable or fixed location encoding approaches (see Fig. 3), while the results are discussed in Section 3.3.2:
Figure 3: Visualisation of what numerical values are encoded through location embeddings using three spatial neighbours (k=3)
@@
  • Rank: We encode the rank of nearest neighbours, i.e. the k in the kNN, with a $512$ dimensional 1D1-D sinusoidal encoding, similarly to the Transformer 1-D positional encoding (Vaswani et al., 2017). Embeddings are concatenated with the text or visual features and merged through a fully connected (FC) layer.
@@
  • Learnable: We encode the rank of nearest neighbours for both text and image embeddings with learnable embeddings, which are then summed with the corresponding vision or text features. The embeddings are initialized with a uniform distribution between $1/h_{dim}^{1/2},1/{h_{dim}^{1/2}}]$.
@@
  • Distance: We encode the Euclidean distance in meters from the origin to the target nearest neighbour with 1D1-D sinusoidal encoding with $h_{dim}=512$ dimensions up to a maximum distance of 10,000m10,000~m. We concatenate the distance embeddings with their corresponding text or visual embeddings and combine them through a FC layer.
  • Coordinates: We encode the geographic coordinates for each location with a 2D2-D sinusoidal multi-scale encoding, similarly to the “grid” approach in Mai et al. (2020). The geographic coordinates are scaled to the range $[0,1]$ based on the min-max coordinates of the study area. We encode each coordinate, i.e. North and East, separately with $h_{dim}/2=256$ dimensions. We concatenate them with their corresponding text or visual embeddings, and finally merge them through a FC layer.
@@
  • Polar: We combine the Euclidean distance from the origin to the target nearest neighbour with the angle between the positive x-axis and the vector pointing from the origin to the nearest neighbour. We use sinusoidal functions to encode both the distance and the angle, similarly to the “polar” approach from Mai et al. (2020). We concatenate the distance, the angle and the features embedding together and combine them through a FC layer.
@@
After encoding the location for all inputs, each text and visual embedding can be considered as an input token without specific order, since the position relative to the origin is preserved.
Figure 4: Illustration of the attention between multimodal tokens in the fusion module, where nodes represent the global classification token (CLS), image tokens $v_{i}$ and text tokens $t_{i}$, and edge width denotes pairwise attention strength.
2.4 Multi-modal fusion with the transformer-based module
@@
Once all inputs have been encoded into a common embedding space, fusion can be performed. The fusion module is composed of a transformer encoder, a layer-normalisation layer, a global pooling layer on the global classification (CLS) token and a projection layer that leads to the final model predictions. The attention mechanisms in the transformer encoder enable the fusion of the multi-modal tokens. The motivation behind this design is that the attention mechanism dynamically weights spatial neighbours (see Fig. 4), emphasising the most informative neighbours, while ignoring distant or irrelevant ones, thereby learning a context-aware spatial representation.
@@ 2.5.1 EcoWikiRS
The EcoWikiRS (Zermatten et al., 2025) dataset contains over $90,000$ text-image pairs, combining high-resolution RGB aerial imagery from Switzerland (50cm50~cm resolution, $100~m\times 100~m$) with sentences from Wikipedia describing the habitat of a species observed at the location of the images. The geolocation of species observations from the Global Biodiversity Information Facility (GBIF) [^1] https://www.gbif.org/ was used to pair species habitat descriptions with aerial images. A specificity of this dataset is that images are associated with more than one sentence, since several sentences are usually extracted from each Wikipedia article. The dataset is split into train, validation and test following a spatial block split, with a block size of roughly 20km20~km, and nearest neighbours are sampled only within the same split. This avoids performance inflation due to spatially autocorrelated training and test samples (Kattenborn et al., 2022).
@@ 2.5.2 SWECO25
We assess the contribution of each modality in the EcoWikiRS dataset by evaluating its capacity to predict environmental variables. For this task, we use the SWECO25 (Külling et al., 2024) dataset, a data cube of more than $5,000$ environmental rasters at 25m25~m spatial resolution covering Switzerland. For each location in EcoWikiRS, we extract average values covering the surface of the input image ($100\times 100~m$) for $103$ selected SWECO25 variables. We standardise the numerical values across the training set with zero mean and unit variance. The $103$ variables are grouped into seven thematic groups, which are described below, while the Supplementary Section 5.3 provides more details on the selection, preprocessing and individual description of the SWECO25 variables:
@@ 2.6 Experimental setup
During training, random token masking is applied on the visual and textual tokens, with the probability of $30\%$, to reduce overfitting and increase model robustness to varying NN distances. We use the habitat sentences from the EcoWikiRS data, i.e. sentences relative to the habitat section of the article only, since they were shown to perform best in Zermatten et al. (2025). We use the four sentences with the highest cosine similarity to the image in the SkyCLIP latent space, i.e., the top-j sentence with $j=4$. If fewer than four sentences are available, sentences are chosen with replacement. If not mentioned otherwise, a polar location encoding is applied. We evaluated the regression of environmental variables with the coefficient of determination ($R^{2}$).
@@ 3.1 Combining text and images in a spatial context
Figure 8: Attention visualisations for the “Text + Images” model on the EcoWikiRS test set. Attention values are log-maximum scores taken across all layers and scaled to $[0,1]$, and they are represented as means and standard deviations, per distance bins of 100m100~m.
Fig. 7 illustrates how the number of nearest neighbours affects the accuracy metrics. Performance increases steadily up to $10$ NN, after which it stagnates or even drops for some models. Including distant neighbours yields no meaningful performance gain while increasing computational cost, arguing for a conservative approach to data integration. This behaviour is becomes clearer when looking at the Suppl. Figs  15-15, which illustrates the mean distance to NN, and the spatial autocorrelation between input observations, measured as the feature similarity as a function of distance. Higher similarity between features occurs up to a distance of 1,000m1,000~m for visual embeddings, and 700m700~m for the text ones, on average, which is close to the average distance of the $10^{th}$ NN, which equals 880m880~m. This suggests that the spatial attention mechanism implicitly captures spatial dependencies and effectively uses them to weight input features.
Fig. 8 confirms this hypothesis and shows the attention for all visual and text tokens as a function of distance, in the test set. The attention score is computed as the logarithm of the maximum attention score across all transformer encoder layers for the best “Text + Images” model and plotted as distance bins of 100m100~m. Both modalities exhibit their highest attention values at short distances, which then monotonically decrease up to a distance of about 1,000m1,000~m. The visual tokens exhibit, overall, a higher attention score compared to the text tokens.
Overall, this supports a conservative yet clever integration of features, while leveraging a few NNs and a single image already achieves a notable gain in performance compared to methods involving a single modality, or without spatial context.
@@ 3.2 Specific contribution of text and images for different types of environmental variables
For all the thematic groups, a gain in performance is obtained when adding spatial context, but to a varying degree. The land use/land cover, bioclimatic, edaphic and population groups are associated with globally high $R^{2}$ scores, and received a significant boost when adding spatial context, with a large relative increase in performance from $0.47$ to $0.61$ ($30\%$) for bioclimatic variables and from $0.40$ to $0.54$ ($35\%$) for the edaphic group. The overall highest $R^{2}$ scores are obtained for the Vegetation group, both with ($0.75$) and without ($0.74$) spatial context, while this small relative improvement when adding spatial context suggests that features at the origin are sufficient to model this group. For the remaining groups, geological and hydrological variables, the model performance remains low in both scenarios, with only a slight increase with spatial context. These factors are discretised from categorical variables and seem hardly predictable from the RS perspective; adding context information only provides a moderate benefit.
@@ 3.3.1 Integrating text
Fig. 11 (a) illustrates the results for the “Text only” setting. More sentences lead to an increase in the $R^{2}$ correlation coefficient from $0.10$ to $0.15$, as the models better capture relevant context, improving the overall results. Selecting a small number of informative sentences (top-$j$) is more effective than simply increasing the number of sentences for a given location. For instance, selecting the top-4 sentences leads to an $R^{2}$ value of $0.17$ compared to $0.13$ when randomly sampling 4 sentences. When incorporating top-$4$ sentences from NN, i.e. a total of 40 sentences, the performance climbs to $0.27$. For the “Text + Images” model in Fig. 11 (b), we do not see a significant increase in performance between selecting top-$j$ sentences or random sentences, probably due to the fact that the performance is driven mostly by the visual modality. However, similar to the “Text only” setting, integrating text from spatial neighbours leads to the overall best performance.
@@ 3.3.2 Integrating location
In the presence of visual neighbours, adding absolute coordinates does not seem to provide any advantage with respect to adding no location information at all (None). As coordinates encoding offers a spatial resolution of approximately 1km1~km when compressed into 512dimensions512~dimensions, it is possibly too coarse for the considered task, with the $10^{th}$ NN lying at 1km1~km on average (Suppl. Fig. 15).
We can see better performances with approaches that encode the relative position of NN to the origin location. The distance performs only slightly better than None, since absolute distance fails to separate neighbours located at similar ranges, but in different directions. Polar and rank obtain the best performance together with learnable. Rank and learnable seem to be effective because they order NN by decreasing order of spatial correlation, thus theoretically of decreasing importance. Polar encoding is more effective at encoding the spatial context than distances: it provides roughly 40m40~m resolution for distances up to 10km10~km, and the angle resolves neighbours located at similar ranges but in different directions. Overall, Polar obtains similar or best performance, but with fewer trainable parameters, motivating our choice for other experiments.
@@ 4 Discussion and Conclusion
Geospatial data integration remains a central challenge in ecological modelling, particularly when combining gridded observations (e.g. remote sensing images) with sparse and irregularly distributed records (e.g. GBIF descriptions, camera trap images). We propose GAMMA to overcome this challenge through explicit location encoding. By representing all inputs as location-aware embeddings and leveraging a transformer encoder to model spatial and cross-modal relationships, GAMMA preserves spatial structure without the need for interpolating sparse observations into continuous spatial fields.
@@
We evaluate GAMMA on the simultaneous prediction of 103 environmental variables from SWECO25 (Külling et al., 2024), integrating aerial imagery and Wikipedia-derived habitat descriptions over the Swiss territory. Our results demonstrate that integrating two modalities, text and RS imagery, outperforms approaches relying only on one individual modality, i.e. “Text only” or “Image only” approach. Moreover, the proposed integration of spatial neighbours with geolocation embeddings consistently provides an additional gain in performance, underlining the benefits of incorporating spatial information for modelling geospatial processes. We analyse the benefits of each modality: Text proves to be most beneficial for variables governed by broad-scale processes (climate, edaphic, population), while variables driven by local features (vegetation) benefit mostly from features extracted from imagery. Variables that were poorly predicted from remote sensing alone (e.g., geological and hydrological variables) exhibit a marginal gain from spatial context, and no clear benefit from text integration, suggesting that insufficient relevant information was found in either modality. This points to a limitation of our experimental setup: performance is ultimately bounded by the information content of the available modalities, and some environmental processes may require dedicated observational sources not considered here. A further limitation concerns the risk of spatial overfitting. Despite the spatial block split, the model may exploit spurious correlations or geographic sampling biases rather than learning transferable semantic content. In particular, it remains unclear whether specific “prototype” words dominate predictions at certain locations and how this affects generalisation (Chen et al., 2023). Disentangling genuine semantic learning from geographic memorisation is an open problem that warrants dedicated investigation in future work.
Overall, this work demonstrates that representing heterogeneous ecological data as location-aware tokens enables flexible and scalable multimodal integration. We hope this work facilitates further integration of various data sources in ecological applications at a large scale.